if the age of a father is 30 years more than his son, after 4 years the age of the father will be four time than his son. what is the age of the son?

Answer by addingup(3677)

f = s+30
f+4 = 4(s+4) substitute for f on left and distribute 4 on right
s+30+4 = 4s+16
s+34 = 4s+16
18 = 3s
3s = 18
s = 6
The son is 6 and the father is 6+30 = 36
In 4 years:
f: 36+4 = 40
s: 6+4 = 10 and 10*4 = 40 Correct
:
John